# Hybrid Memetic Search for Electric Vehicle Routing with Time Windows, Simultaneous Pickup-Delivery, and Partial Recharges

### v1

@@ V-A2 Parameter Setting
We determined the parameter settings of HMA through preliminary experiments. The settings for different problem scales are detailed in Table I. Specifically, HMA requires setting six parameters: $G$, $N$, $\alpha$, $B$, $sr$, and $w_{1},w_{2}$. $G$ denotes the longest consecutive generation without improvement in HMA, and $N$ is the population size in MS (see Algorithm 3). $\alpha$ and $B$ refer to the proportionality parameter and the maximum number of generations in PSI, respectively. $sr$ specifies the selection range in station ranking list (see Algorithm 1). $w_{1}$ and $w_{2}$ are the lower and upper bounds for the proportion of customers removed in the destroy-repair operator (see Algorithm 3). All the experiments were run on an Intel Xeon Gold 6240 machine with 377 GB RAM and 32 cores (2.60 GHz, 25 MB Cache), running Rocky Linux 8.5. HMA was implemented in C++.
@@
Instance m best avg time m best avg time m best avg time
c101 12 13043.40 13043.42 270.56 12 13057.80 13063.54 205.53 12 13043.38 13043.43 12.57
c102 11 12056.80 12920.23 393.95 11 12073.10 12944.34 329.34 10 11089.80 12033.33 83.87
c103 11 12004.70 12026.90 317.63 10 11134.90 11917.80 504.47 10 10945.45 10945.48 136.32
c104 10 10872.80 11353.78 442.55 10 10870.70 10876.49 427.43 9 9890.98 10767.31 231.74
c105 11 12023.80 12341.60 394.88 11 12034.10 12068.86 409.38 11 12030.81 12130.20 64.75
c106 11 12013.10 12438.06 458.04 11 12025.70 12059.29 305.45 10 11040.32 11825.37 73.50
c107 11 12006.40 12023.97 378.24 11 12026.70 12046.38 276.09 10 11021.16 11815.55 67.93
c108 11 11994.70 12016.10 407.11 10 11025.80 11822.60 391.00 10 11006.10 11597.49 137.57
c109 10 11042.20 11885.30 502.22 10 10941.00 11180.77 524.19 10 10911.47 10936.41 184.63
c201 4 4629.95 4629.95 26.41 4 4678.37 4703.43 274.65 4 4629.95 4629.95 9.55
c202 4 4629.95 4629.95 192.19 4 4664.26 4706.94 277.38 4 4629.95 4629.95 34.02
c203 4 4632.27 4690.06 520.20 4 4641.45 4734.31 349.57 4 4629.95 4629.95 47.47
c204 4 4633.08 4665.78 563.25 4 4660.64 4737.07 503.66 4 4628.91 4628.91 102.73
c205 4 4629.95 4629.95 54.00 4 4629.95 4629.95 88.12 4 4629.95 4629.95 31.58
c206 4 4629.95 4629.95 149.83 4 4629.95 4629.95 142.64 4 4629.95 4630.26 39.25
c207 4 4629.95 4629.95 179.74 4 4629.95 4635.27 182.89 4 4629.95 4630.89 62.34
c208 4 4629.95 4629.95 200.06 4 4629.95 4629.95 183.86 4 4629.95 4629.95 50.89
r101 18 19633.80 19939.79 459.20 18 19640.60 19661.15 476.75 17 18636.96 19512.43 75.46
r102 17 18470.80 19292.16 496.96 16 17474.10 17696.35 561.10 16 17424.75 17526.29 115.80
r103 15 16296.50 17050.75 499.65 14 15280.30 15306.17 449.49 13 14209.21 14721.75 187.46
r104 13 14141.10 14255.53 433.34 12 13084.30 13111.31 538.23 11 12051.19 12066.53 285.75
r105 15 16389.20 17212.83 477.88 14 15471.30 16346.10 429.92 14 15366.33 16159.23 146.98
r106 15 16292.00 16836.67 492.99 14 15314.80 15441.68 524.23 13 14265.55 14373.31 222.14
r107 13 14168.90 15016.67 478.40 12 13140.10 13669.50 550.64 12 13112.66 13118.04 226.39
r108 12 13079.80 13531.30 468.60 11 12073.70 12998.17 521.61 11 12016.74 12025.10 289.00
r109 14 15237.30 15674.51 533.66 13 14220.80 14468.12 523.07 12 13203.74 13993.94 237.95
r110 13 14170.20 14905.73 370.99 12 13114.30 13544.76 541.43 11 12070.04 12482.82 270.95
r111 12 13144.20 14584.19 489.45 12 13148.80 13965.98 503.08 12 13070.64 13077.15 225.63
r112 12 13155.60 14053.56 331.29 12 13044.10 13078.65 597.35 11 12001.79 12012.62 233.83
r201 4 5192.33 5216.92 506.12 4 5276.75 5363.04 131.64 3 4268.88 4985.55 264.02
r202 3 4250.70 5020.88 483.79 3 4193.33 4940.22 615.58 3 4054.75 4063.20 363.02
r203 3 3942.74 4352.52 609.82 3 3985.02 4060.18 577.82 3 3898.68 3910.77 541.38
r204 3 3820.72 3854.31 553.01 3 3793.76 3827.81 615.67 2 2815.58 3544.13 438.26
r205 3 4055.28 4124.64 514.20 3 4065.06 4126.45 253.23 3 3995.35 4003.09 512.99
r206 3 3978.10 4065.05 531.38 3 3991.44 4047.16 550.82 3 3925.95 3938.12 600.35
r207 3 3878.91 3910.07 463.55 3 3881.97 3918.29 616.95 2 2851.75 3534.37 494.20
r208 3 3791.27 3829.39 596.95 3 3732.80 3776.20 629.30 2 2740.06 2756.80 589.42
r209 3 3975.64 4015.90 467.75 3 3933.55 3977.24 538.01 3 3877.24 3885.71 636.14
r210 3 3920.37 3984.78 530.97 3 3926.79 3961.42 520.11 3 3846.62 3854.33 555.78
r211 3 3814.42 3893.38 580.24 3 3824.47 3857.62 561.62 3 3772.67 3789.62 477.61
rc101 16 17667.70 18513.67 504.48 16 17696.20 17741.63 442.05 15 16661.53 17462.90 152.22
rc102 16 17576.80 17909.78 392.20 15 16558.20 16628.46 422.70 14 15510.31 15618.29 256.61
rc103 14 15366.90 16245.06 537.04 13 14358.20 14999.16 485.96 12 13343.56 13648.08 297.57
rc104 13 14270.50 14315.17 447.54 12 13222.50 13261.65 490.66 11 12150.47 12168.35 388.45
rc105 15 16500.90 16933.42 458.25 14 15470.70 15820.90 449.50 14 15447.52 15454.71 164.64
rc106 14 15432.50 16069.05 444.10 13 14448.30 15249.17 406.17 13 14388.12 14609.85 190.86
rc107 13 14313.40 14437.31 540.67 12 13276.50 13386.51 518.89 11 12234.76 13048.02 232.42
rc108 12 13226.00 13891.71 435.95 12 13184.70 13214.50 504.05 11 12140.28 12152.42 271.54
rc201 4 5504.77 5819.06 494.52 4 5617.75 5786.48 226.81 4 5444.90 5450.77 254.36
rc202 4 5324.64 5442.41 416.84 4 5436.63 5541.80 138.34 4 5240.40 5249.87 192.66
rc203 4 5109.88 5177.69 452.57 4 5086.44 5159.15 531.94 3 4078.55 4100.32 597.82
rc204 3 4036.49 4525.03 523.68 3 3962.88 4252.28 631.65 3 3892.97 3902.06 528.96
rc205 4 5260.14 5338.50 426.74 4 5285.41 5375.54 221.24 4 5149.13 5152.79 368.18
rc206 4 5234.55 5289.90 470.82 4 5210.57 5275.72 395.42 3 4200.69 4323.33 517.29
rc207 3 4150.60 4930.81 488.10 3 4197.81 4650.01 607.24 3 3979.65 4019.57 490.17
rc208 3 3977.50 4046.09 535.72 3 3920.17 4002.79 527.17 3 3838.78 3857.49 457.65
avg 8.48 9568.86 9905.20 435.54 8.18 9266.06 9479.93 432.20 7.77 8807.09 9030.14 262.90
V-B Results on Small-scale and Medium-scale Instances
We compared HMA with two recently proposed algorithms, Adapt-CMSA-STD and Adapt-CMSA-SETCOV, on the akb set. Since their source code is not publicly available, we obtained the testing results directly from the original publications [11, 17]. To ensure a fair comparison, we rescaled the computation time to a common measure considering the CPUs used. Specifically, we rescaled the times based on the PassMark score of the CPUs used in the studies relative to our 2.60 GHz Intel Xeon Gold 6240. The PassMark score, as the performance of a single core of the respective CPU, for the 2.93 GHz X5670 used in [11, 17] is 1398, while the score of our CPU is 1990. The rescaled time in seconds is given in Tables II and III. Additionally, the run time limit on the small-scale and medium-scale instances is rescaled to 105 and 630 seconds.
Tables II and III are structured as follows. The first column lists the instance names. The “m” columns indicate the number of vehicles used in the respective solutions. For Adapt-CMSA-STD, Adapt-CMSA-SETCOV, and HMA, the “best” column refers to the $TC$ of the best solution found in 10 independent runs, while the “avg’ column refers to the average $TC$ of the 10 runs. Finally, the “time” column shows the computation time (in seconds) for CPLEX and the average computation time for Adapt-CMSA-STD, Adapt-CMSA-SETCOV, and HMA to find the best solutions in each run, respectively. For each instance, the best performance w.r.t. “best” and “avg” are highlighted in gray, respectively.
@@ V-B Results on Small-scale and Medium-scale Instances
HMA w/o BCD HMA
Instance m best avg±std time m best avg±std time p-value
jd200_1 44 71043.42 72243.96±614.67 1884.42 44 70979.75 72348.31±724.19 1821.75 0.6501
jd200_2 39 59660.84 60785.77±551.51 1892.70 39 59460.27 60688.60±697.48 1825.64 0.7624
jd200_3 44 66634.23 67043.56±269.90 1872.56 43 66254.27 66999.61±526.33 1839.70 0.8798
jd200_4 39 60988.30 62033.06±488.03 1873.78 39 60941.60 61753.23±523.17 1843.03 0.2568
jd400_1 79 116029.37 117199.07±954.98 5784.07 79 115981.66 116987.12±952.18 5443.33 0.4963
jd400_2 77 116527.33 117584.91±736.78 5734.43 77 116205.05 117098.09±563.09 5430.44 0.1041
jd400_3 76 109030.69 109934.10±508.53 5576.75 76 108441.53 109140.65±451.85 5421.66 0.0032
jd400_4 83 128988.11 129937.62±710.96 5761.35 83 128894.04 129654.11±858.13 5428.58 0.1509
jd600_1 112 167369.71 169339.19±1098.95 9351.29 112 167297.70 168836.18±1087.42 9033.06 0.3258
jd600_2 113 171996.76 174044.48±1610.00 9643.47 113 171512.97 173480.89±1384.65 9019.82 0.3447
jd600_3 112 154336.98 155381.50±773.48 9610.60 111 153092.64 154325.16±840.34 9037.23 0.0191
jd600_4 104 150513.45 151553.99±649.75 9753.66 103 150123.89 150975.72±609.08 9049.35 0.0343
jd800_1 147 216364.67 217948.77±1027.41 13710.86 146 214710.04 216810.78±1542.74 12625.32 0.0757
jd800_2 148 226218.76 228162.86±1987.21 13523.34 149 223610.37 226527.29±2364.16 12619.90 0.0963
jd800_3 148 206772.88 208548.00±1067.70 13401.35 147 205215.92 207375.35±1432.89 12619.30 0.0494
jd800_4 136 196076.58 197143.07±637.84 13603.29 134 194004.95 196257.28±1325.29 12621.05 0.0696
jd1000_1 179 261338.90 262810.90±1082.93 17483.69 178 257410.81 260877.67±2210.64 16224.57 0.0343
jd1000_2 183 276597.69 279318.78±1341.59 16643.43 184 274079.42 276515.30±1881.23 16224.80 0.0028
jd1000_3 184 256229.79 258040.35±1467.13 16631.34 182 252772.63 256102.57±2369.95 16221.87 0.0494
jd1000_4 169 243354.91 244808.00±1235.27 17552.66 168 240559.58 242952.46±2044.65 16218.06 0.0284
Based on the numerical results in Table II, the following observations can be made. On small-scale instances, CPLEX, Adapt-CMSA-STD, Adapt-CMSA-SETCOV, and HMA are comparable, showing no significant difference in their best or average values due to the small solution space. Moreover, it is worth mentioning that HMA exhibits minimal computation time variation as the problem scale increases, with an average computation time of 2.20 seconds, which is much shorter than that of the other three algorithms.
Generally, the performance of algorithms on instances of larger scales is of more interest. Overall, HMA is the best-performing algorithm on medium-scale instances as shown in Table III. In the $\sim$ and $\sim$ subsets, the best and average solution quality obtained by HMA are consistently better than those of Adapt-CMSA-STD and Adapt-CMSA-SETCOV. Note that for the akb set, minimizing the number of the used EVs has a higher priority than minimizing $TD$. It is worth mentioning that HMA obtained competitive solutions with fewer EVs in 23 out of 56 instances. More specifically, these instances include 4 out of 9 in the c1$\sim$ subset, 7 out of 12 in the r1$\sim$ subset, 4 out of 11 in the r2$\sim$ subset, 6 out of 8 in the rc1$\sim$ subset, and 2 out of 8 in the rc2$\sim$ subset. Such results indicate that HMA performed both strongly and robustly. Moreover, HMA required significantly less computation time, taking only about 60% of the time compared to the other two algorithms. In summary, the observations demonstrate that HMA outperforms all recent EVRP-TW-SPD algorithms across a broad range of problem instances, and in particular, it has significantly raised the performance bar on the akb set.

### v2

@@ V-A2 Parameter Setting
We determined the parameter settings of HMA through preliminary experiments. The settings for different problem scales are detailed in Table I. Specifically, HMA requires setting seven parameters: $G_{1}$, $G_{2}$, $N$, $\alpha$, $B$, $sr$, and $w_{1},w_{2}$. $G_{1}$ and $G_{2}$ denote the longest consecutive iterations/generations without improvement in LNS and MS of HMA, respectively, while $N$ represents the population size in MS (see Algorithm 3). $\alpha$ and $B$ refer to the proportionality parameter and the maximum number of generations in PSI, respectively. $sr$ specifies the selection range in station ranking list (see Algorithm 1). $w_{1}$ and $w_{2}$ are the lower and upper bounds for the proportion of customers removed in the destroy-repair operator (see Algorithm 3). All the experiments were run on an Intel Xeon Gold 6240 machine with 377 GB RAM and 32 cores (2.60 GHz, 25 MB Cache), running Rocky Linux 8.5. HMA was implemented in C++.
V-B Results on Small-scale and Medium-scale Instances
We compared HMA with two recently proposed state-of-the-art algorithms, Adapt-CMSA-STD and Adapt-CMSA-SETCOV, on the akb set. Since their source code is not publicly available, we obtained the testing results directly from the original publications [13, 19]. To ensure a fair comparison, we rescaled the computation time to a common measure considering the CPUs used. Specifically, we rescaled the time based on the PassMark score of the CPUs used in the studies relative to our 2.60 GHz Intel Xeon Gold 6240. The PassMark score, which reflects the performance of a single CPU core, is 1398 for the 2.93 GHz X5670 used in[13, 19], whereas our CPU scores 1990. The rescaled times (in seconds) are reported in Tables II and III. Additionally, after rescaling, the run time limits for the small-scale and medium-scale instances were set to 105 and 630 seconds, respectively.
Tables II and III are structured as follows. The first column lists the instance names. The “m” columns indicate the number of EVs used in the respective solutions. For Adapt-CMSA-STD, Adapt-CMSA-SETCOV, and HMA, the “best” column refers to the $TC$ of the best solution found in 10 independent runs, while the “avg” column refers to the average $TC$ of the 10 runs. The “time” column shows the computation time (in seconds) for CPLEX and the average computation time for Adapt-CMSA-STD, Adapt-CMSA-SETCOV, and HMA to find the best solutions in each run, respectively. Finally, the “gap” column shows the relative improvement of the best solution obtained by HMA compared to the existing best-known solution. Specifically, $gap=[TC_{\text{HMA}}-TC_{\text{best known}}]/TC_{\text{best known}}$. A negative value of $gap$ indicates that HMA has achieved an improvement in solution quality. For each instance, the best “best” value and the best “avg” value are highlighted in gray.
Based on the numerical results in Table II, the following observations can be made. On small-scale instances, the three metaheuristic algorithms were able to obtain optimal solutions in much shorter computation time compared to CPLEX, and even found better solutions than CPLEX in instances r202C15 and rc204C15. Overall, Adapt-CMSA-STD, Adapt-CMSA-SETCOV, and HMA exhibit comparable performance, with no statistically significant differences in their best or average solution quality, likely due to the small solution space of these instances (see Section V-D for the statistical test). The $gap$ value is $0.00\%$ for nearly all instances except r202C15, indicating that HMA achieved solution quality equivalent to the best-known solutions in most cases. Furthermore, HMA demonstrates minimal variation in computation time as the problem scale increases, with an average runtime of 2.20 seconds, which is much shorter than that of the other two metaheuristic algorithms.
Generally, the performance of algorithms on instances of larger scales is of more interest. Overall, HMA is the best-performing algorithm on medium-scale instances as shown in Table III. In the $\sim$ and $\sim$ subsets, the best and average solution quality obtained by HMA are consistently better than those of Adapt-CMSA-STD and Adapt-CMSA-SETCOV. Moreover, in the $\sim$ subsets, HMA outperformed the other two algorithms in all except four instances. Note that for the akb set, minimizing the number of used EVs has a higher priority than minimizing $TD$, as indicated by the settings $\mu_{1}=1000$ and $\mu_{2}=1$ in Eq. (5a). It is worth mentioning that HMA obtained competitive solutions with fewer EVs in 23 out of 56 medium-scale instances. More specifically, these instances include 4 out of 9 in the c1$\sim$ subset, 7 out of 12 in the r1$\sim$ subset, 4 out of 11 in the r2$\sim$ subset, 6 out of 8 in the rc1$\sim$ subset, and 2 out of 8 in the rc2$\sim$ subset. It is observed that in all the aforementioned instances, the $gap$ values are below $-5.00\%$, indicating a significant improvement. For the remaining instances where the same number of EVs is used, the $gap$ values are nearly $0.00\%$, yet remain negative, which suggests that HMA still achieved solution quality better than the best-known solutions in those cases. Considering that customer distribution and constraint intensity vary across different instance types (see Section V-A1), such results indicate that HMA performed both strongly and robustly. Moreover, HMA required considerably less computation time, using only about 60% of the time consumed by the other two algorithms.
Compared to Adapt-CMSA-STD and Adapt-CMSA-SETCOV, HMA exhibits a remarkable performance advantage on medium-scale instances. This improvement is primarily due to the fact that, unlike Adapt-CMSA-STD and Adapt-CMSA-SETCOV, which maintain and refine only a single elite solution throughout the search process, HMA employs a memetic search that initializes the population based on the elite solution obtained from the large neighborhood search. This strategy enables effective exploitation of the high-quality information contained in the elite solution, while maintaining population diversity to reduce the risk of premature convergence. In addition, the combined use of PSI and SSI, helps avoid getting trapped in poor-quality local optima. The effectiveness of these novel components integrated into HMA will be further analyzed in Section V-E.
In summary, the observations presented above collectively indicate that HMA outperforms recent state-of-the-art algorithms across a broad range of problem instances. Notably, it has significantly raised the performance bar on the akb set.
@@ V-B Results on Small-scale and Medium-scale Instances
Instance name m best avg time m best avg time m best avg time gap
c101 12 13043.40 13043.42 270.56 12 13057.80 13063.54 205.53 12 13043.38 13043.43 12.57 0.00%
c102 11 12056.80 12920.23 393.95 11 12073.10 12944.34 329.34 10 11089.80 12033.33 83.87 -8.02%
c103 11 12004.70 12026.90 317.63 10 11134.90 11917.80 504.47 10 10945.45 10945.48 136.32 -1.70%
c104 10 10872.80 11353.78 442.55 10 10870.70 10876.49 427.43 9 9890.98 10767.31 231.74 -9.01%
c105 11 12023.80 12341.60 394.88 11 12034.10 12068.86 409.38 11 12030.81 12130.20 64.75 0.06%
c106 11 12013.10 12438.06 458.04 11 12025.70 12059.29 305.45 10 11040.32 11825.37 73.50 -8.10%
c107 11 12006.40 12023.97 378.24 11 12026.70 12046.38 276.09 10 11021.16 11815.55 67.93 -8.21%
c108 11 11994.70 12016.10 407.11 10 11025.80 11822.60 391.00 10 11006.10 11597.49 137.57 -0.18%
c109 10 11042.20 11885.30 502.22 10 10941.00 11180.77 524.19 10 10911.47 10936.41 184.63 -0.27%
c201 4 4629.95 4629.95 26.41 4 4678.37 4703.43 274.65 4 4629.95 4629.95 9.55 0.00%
c202 4 4629.95 4629.95 192.19 4 4664.26 4706.94 277.38 4 4629.95 4629.95 34.02 0.00%
c203 4 4632.27 4690.06 520.20 4 4641.45 4734.31 349.57 4 4629.95 4629.95 47.47 -0.05%
c204 4 4633.08 4665.78 563.25 4 4660.64 4737.07 503.66 4 4628.91 4628.91 102.73 -0.09%
c205 4 4629.95 4629.95 54.00 4 4629.95 4629.95 88.12 4 4629.95 4629.95 31.58 0.00%
c206 4 4629.95 4629.95 149.83 4 4629.95 4629.95 142.64 4 4629.95 4630.26 39.25 0.00%
c207 4 4629.95 4629.95 179.74 4 4629.95 4635.27 182.89 4 4629.95 4630.89 62.34 0.00%
c208 4 4629.95 4629.95 200.06 4 4629.95 4629.95 183.86 4 4629.95 4629.95 50.89 0.00%
r101 18 19633.80 19939.79 459.20 18 19640.60 19661.15 476.75 17 18636.96 19512.43 75.46 -5.08%
r102 17 18470.80 19292.16 496.96 16 17474.10 17696.35 561.10 16 17424.75 17526.29 115.80 -0.28%
r103 15 16296.50 17050.75 499.65 14 15280.30 15306.17 449.49 13 14209.21 14721.75 187.46 -7.01%
r104 13 14141.10 14255.53 433.34 12 13084.30 13111.31 538.23 11 12051.19 12066.53 285.75 -7.90%
r105 15 16389.20 17212.83 477.88 14 15471.30 16346.10 429.92 14 15366.33 16159.23 146.98 -0.68%
r106 15 16292.00 16836.67 492.99 14 15314.80 15441.68 524.23 13 14265.55 14373.31 222.14 -6.85%
r107 13 14168.90 15016.67 478.40 12 13140.10 13669.50 550.64 12 13112.66 13118.04 226.39 -0.21%
r108 12 13079.80 13531.30 468.60 11 12073.70 12998.17 521.61 11 12016.74 12025.10 289.00 -0.47%
r109 14 15237.30 15674.51 533.66 13 14220.80 14468.12 523.07 12 13203.74 13993.94 237.95 -7.15%
r110 13 14170.20 14905.73 370.99 12 13114.30 13544.76 541.43 11 12070.04 12482.82 270.95 -7.96%
r111 12 13144.20 14584.19 489.45 12 13148.80 13965.98 503.08 12 13070.64 13077.15 225.63 -0.56%
r112 12 13155.60 14053.56 331.29 12 13044.10 13078.65 597.35 11 12001.79 12012.62 233.83 -7.99%
r201 4 5192.33 5216.92 506.12 4 5276.75 5363.04 131.64 3 4268.88 4985.55 264.02 -17.78%
r202 3 4250.70 5020.88 483.79 3 4193.33 4940.22 615.58 3 4054.75 4063.20 363.02 -3.30%
r203 3 3942.74 4352.52 609.82 3 3985.02 4060.18 577.82 3 3898.68 3910.77 541.38 -1.12%
r204 3 3820.72 3854.31 553.01 3 3793.76 3827.81 615.67 2 2815.58 3544.13 438.26 -25.78%
r205 3 4055.28 4124.64 514.20 3 4065.06 4126.45 253.23 3 3995.35 4003.09 512.99 -1.48%
r206 3 3978.10 4065.05 531.38 3 3991.44 4047.16 550.82 3 3925.95 3938.12 600.35 -1.31%
r207 3 3878.91 3910.07 463.55 3 3881.97 3918.29 616.95 2 2851.75 3534.37 494.20 -26.48%
r208 3 3791.27 3829.39 596.95 3 3732.80 3776.20 629.30 2 2740.06 2756.80 589.42 -26.60%
r209 3 3975.64 4015.90 467.75 3 3933.55 3977.24 538.01 3 3877.24 3885.71 636.14 -1.43%
r210 3 3920.37 3984.78 530.97 3 3926.79 3961.42 520.11 3 3846.62 3854.33 555.78 -1.88%
r211 3 3814.42 3893.38 580.24 3 3824.47 3857.62 561.62 3 3772.67 3789.62 477.61 -1.09%
rc101 16 17667.70 18513.67 504.48 16 17696.20 17741.63 442.05 15 16661.53 17462.90 152.22 -5.69%
rc102 16 17576.80 17909.78 392.20 15 16558.20 16628.46 422.70 14 15510.31 15618.29 256.61 -6.33%
rc103 14 15366.90 16245.06 537.04 13 14358.20 14999.16 485.96 12 13343.56 13648.08 297.57 -7.07%
rc104 13 14270.50 14315.17 447.54 12 13222.50 13261.65 490.66 11 12150.47 12168.35 388.45 -8.11%
rc105 15 16500.90 16933.42 458.25 14 15470.70 15820.90 449.50 14 15447.52 15454.71 164.64 -0.15%
rc106 14 15432.50 16069.05 444.10 13 14448.30 15249.17 406.17 13 14388.12 14609.85 190.86 -0.42%
rc107 13 14313.40 14437.31 540.67 12 13276.50 13386.51 518.89 11 12234.76 13048.02 232.42 -7.85%
rc108 12 13226.00 13891.71 435.95 12 13184.70 13214.50 504.05 11 12140.28 12152.42 271.54 -7.92%
rc201 4 5504.77 5819.06 494.52 4 5617.75 5786.48 226.81 4 5444.90 5450.77 254.36 -1.09%
rc202 4 5324.64 5442.41 416.84 4 5436.63 5541.80 138.34 4 5240.40 5249.87 192.66 -1.58%
rc203 4 5109.88 5177.69 452.57 4 5086.44 5159.15 531.94 3 4078.55 4100.32 597.82 -19.82%
rc204 3 4036.49 4525.03 523.68 3 3962.88 4252.28 631.65 3 3892.97 3902.06 528.96 -1.76%
rc205 4 5260.14 5338.50 426.74 4 5285.41 5375.54 221.24 4 5149.13 5152.79 368.18 -2.11%
rc206 4 5234.55 5289.90 470.82 4 5210.57 5275.72 395.42 3 4200.69 4323.33 517.29 -19.38%
rc207 3 4150.60 4930.81 488.10 3 4197.81 4650.01 607.24 3 3979.65 4019.57 490.17 -4.12%
rc208 3 3977.50 4046.09 535.72 3 3920.17 4002.79 527.17 3 3838.78 3857.49 457.65 -2.08%
avg 8.48 9568.86 9905.20 435.54 8.18 9266.06 9479.93 432.20 7.77 8807.09 9030.14 262.90 -5.20%
@@
Adapt-CMSA-STD Adapt-CMSA-SETCOV HMA w/o BCD HMA
Instance name best avg±$\pm$±std time $\Delta\%$ best avg±$\pm$±std time $\Delta\%$ best avg±$\pm$±std time $\Delta\%$ best avg±$\pm$±std time p-value
jd200_1 74060.10 75718.31$\pm$908.58 1782.33 76125.24 78311.24$\pm$1416.76 1803.16 71043.42 72243.96$\pm$614.67 1884.42 70979.75 72348.31$\pm$724.19 1821.75 0.6501
jd200_2 66255.36 67505.53$\pm$655.09 1797.81 71075.99 71965.57$\pm$704.99 1801.42 59660.84 60785.77$\pm$551.51 1892.70 59460.27 60688.60$\pm$697.48 1825.64 0.7624
jd200_3 68219.71 69710.26$\pm$763.65 1686.48 70656.63 71918.09$\pm$792.00 1803.03 66634.23 67043.56$\pm$269.90 1872.56 66254.27 66999.61$\pm$526.33 1839.70 0.8798
jd200_4 72557.01 74701.11$\pm$1482.53 1800.69 9.45% 81696.93 84613.90$\pm$2035.06 1802.73 16.87% 60988.30 62033.06$\pm$488.03 1873.78 0.27% 60941.60 61753.23$\pm$523.17 1843.03 0.2568
jd400_1 141561.75 142694.05$\pm$1128.17 5406.36 143045.76 145872.43$\pm$1772.19 5406.09 116029.37 117199.07$\pm$954.98 5784.07 115981.66 116987.12$\pm$952.18 5443.33 0.4963
jd400_2 136842.18 139007.05$\pm$1738.07 5421.45 137495.23 141242.66$\pm$2324.07 5404.31 116527.33 117584.91$\pm$736.78 5734.43 116205.05 117098.09$\pm$563.09 5430.44 0.1041
jd400_3 118665.20 120745.11$\pm$1413.70 5410.98 120888.17 122891.09$\pm$1530.63 5405.72 109030.69 109934.10$\pm$508.53 5576.75 108441.53 109140.65$\pm$451.85 5421.66 0.0032
jd400_4 137312.31 141331.04$\pm$2085.47 5404.06 13.94% 137863.84 142086.80$\pm$1877.16 5406.27 15.02% 128988.11 129937.62$\pm$710.96 5761.35 0.23% 128894.04 129654.11$\pm$858.13 5428.58 0.1509
jd600_1 213824.87 221616.09$\pm$7471.96 9097.22 210722.03 215743.58$\pm$2767.03 8988.48 167369.71 169339.19$\pm$1098.95 9351.29 167297.70 168836.18$\pm$1087.42 9033.06 0.3258
jd600_2 206849.40 215497.97$\pm$6446.15 9103.28 205943.56 210498.81$\pm$2071.06 9008.39 171996.76 174044.48$\pm$1610.00 9643.47 171512.97 173480.89$\pm$1384.65 9019.82 0.3447
jd600_3 173412.99 176186.74$\pm$1944.15 9243.68 172977.87 175893.80$\pm$1829.45 8887.64 154336.98 155381.50$\pm$773.48 9610.60 153092.64 154325.16$\pm$840.34 9037.23 0.0191
jd600_4 201123.62 211620.07$\pm$7977.66 9009.64 23.91% 197929.76 201306.90$\pm$2408.88 9008.28 22.72% 150513.45 151553.99$\pm$649.75 9753.66 0.35% 150123.89 150975.72$\pm$609.08 9049.35 0.0343
jd800_1 275924.88 313298.97$\pm$20935.22 12617.93 271098.99 276999.84$\pm$4688.42 12557.60 216364.67 217948.77$\pm$1027.41 13710.86 214710.04 216810.78$\pm$1542.74 12625.32 0.0757
jd800_2 269445.25 312947.07$\pm$22210.55 12667.52 268047.17 275025.17$\pm$8658.96 12607.02 226218.76 228162.86$\pm$1987.21 13523.34 223610.37 226527.29$\pm$2364.16 12619.90 0.0963
jd800_3 234915.71 257492.90$\pm$17983.73 13053.00 232052.94 239688.51$\pm$9680.09 12611.89 206772.88 208548.00$\pm$1067.70 13401.35 205215.92 207375.35$\pm$1432.89 12619.30 0.0494
jd800_4 282966.34 297859.86$\pm$13680.91 12613.94 27.33% 260642.67 266243.72$\pm$4288.39 12610.64 23.39% 196076.58 197143.07$\pm$637.84 13603.29 0.94% 194004.95 196257.28$\pm$1325.29 12621.05 0.0696
jd1000_1 378549.15 409496.27$\pm$20008.55 16265.94 327562.85 345987.18$\pm$20488.34 16214.39 261338.90 262810.90$\pm$1082.93 17483.69 257410.81 260877.67$\pm$2210.64 16224.57 0.0343
jd1000_2 369338.23 408955.96$\pm$26354.10 16232.57 335504.67 351073.09$\pm$20290.60 16214.16 276597.69 279318.78$\pm$1341.59 16643.43 274079.42 276515.30$\pm$1881.23 16224.80 0.0028
jd1000_3 314790.77 361696.11$\pm$32886.34 16233.94 286785.82 304623.25$\pm$21348.50 16218.22 256229.79 258040.35$\pm$1467.13 16631.34 252772.63 256102.57$\pm$2369.95 16221.87 0.0494
jd1000_4 349809.78 391564.48$\pm$26270.35 16229.30 37.94% 320782.78 333245.22$\pm$15639.71 16221.33 24.12% 243354.91 244808.00$\pm$1235.27 17552.66 1.24% 240559.58 242952.46$\pm$2044.65 16218.06 0.0284